• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name paschke

    Log entry time 00:11:14 on August19,2009

    Entry number 284932

    keyword=testing left arm switch

    Used new switch on detector signals for L-HRS.  Switch checks out on the bench. 
     Used patch panel connections 10 and 11 for the switching voltage connections.  
    This is a latching switch: apply 12 Volts and you'll toggle the state of the 
    switch.  I think that 10 turns on the parity DAQ output, but I'm not sure. 
    The detector HV (on cable "Gas Cerenkov 10") is Left slot 2, channel 9.
    Run 11427, to test switch... we went from turning on "10" to turning on "11", 
    but found no evidence of signal in either case.  
    Note, the switch takes the detector signal to "socket", and the long cable we 
    made today inverts (goes socket->socket, pin->pin), so the PMT signal should be 
    properly presented on the pin to the ADC.
    Also from run 11427: detector widths are about 2.5 channel difference (left 
    arm) and 4.5 channel difference (right arm).  There is no evidence that either 
    is working, of course.  These are in the 18 bit adcs.  The blumi widths are 
    comparable.  I didn't fix the .db file for blumi8 which is in adc29 but with no 
    dacnoise correction.